iuytyuioiuytyuiop:

It costs ABC electronics company $2.50 per unit to produce a part used in a popular brand of desktop computers. The company has monthly operating expenses of $350 for utilities and $3300 for salaries. Write the linear equation to model the company's monthly expenses in the form of y=mx+b.

Extrinix:

monthly operating bill of \(\color{lime}{$350}\) for utilities, and \(\color{lime}{$3300}\) for salaries, giving us \(\color{lime}{$3650}\) in total, that can be our \(\color{lime}{b}\), then to find \(\color{red}{m}\), \(\color{red}{m}\) should be 2.5 which is \(\color{red}{5/2}\) that is our \(\color{red}{m}\), that give you your answer.

iuytyuioiuytyuiop:

so b * m?

Extrinix:

no, in y=mx+b format it would be \(\color{gray}{y=5/2x+3650}\)

iuytyuioiuytyuiop:

so I add

Extrinix:

no your answer is y=5/2x+3650

iuytyuioiuytyuiop:

thanks